OpenStudy (anonymous):

A potato is put into an oven that has been heated to 350 degrees Fahrenheit. Its temperature as a function of time is give by T(t)=a(1-e^(-kt))+b. The potato was 50 degrees F when it was first put into the oven. If the potato is 60 degrees F after 2 minutes, what is the value of k? Explain

OpenStudy (anonymous):

\work with the difference in the temperatures the initial difference is \(350-50=300\) the difference after 2 minutes is \(350-60=290\)

OpenStudy (anonymous):

this difference in temperatures will decay to zero, so the model for the differences will be something like \(300e^{kt}\) you know if \(t=2\) you get \(290\) so you can find \(k\) by setting \[300e^{2k}=290\] and solve for \(k\)

OpenStudy (anonymous):

\[e^{2k}=\frac{29}{30}\] \[2k=\ln(\frac{2}{30})\] \[k=\frac{\ln(\frac{29}{30})}{2}\]